# On the Limit of Language Models as Planning Formalizers

### v2

@@ 4.3 Models
For both of the LLM-as-planner and LLM-as-formalizer approach, we consider a number of models, including open-source and closed-source LLMs varying in size, including gemma-2-9b|27b-it Team et al. (2024), llama-3.1-8B|70B|405B-Instruct Dubey et al. (2024), DeepSeek-R1-Distill-Llama-8B|70B Guo et al. (2025) [^4] Due to the cost and zero solvability on our easiest dataset, results of Llama-405B and DeepSeek-8B—70B are ommitted., gpt-4o-mini-2024-07-18, gpt-4o-2024-08-06, and o3-mini-2025-01-31555platform.openai.com/docs/models. We query these models using KANI Zhu et al. (2023) with default hyper-parameters. The open-source models are run using 4 RTX A6000 GPUs, averaging about 1062 input and output tokens for the LLM-as-formalizer approach in BlocksWorld-100. To emulate real-life application with minimal user interference, we use zero-shot prompts for all naturalness levels across all datasets (see prompts in Appendix B).
@@ 5.1 Can LLMs formalize?
We seek to understand the extent to which LLMs can act as a formalizer to generate entire PDDL, instead of partial components in previous work. Figure 3 displays the results on our experiments using the most natural sounding descriptions on Heavily Templated BlocksWorld-100, Logistics-100 and MysteryBlocksWorld-100. Results on the most complex domain Barman-100 is omitted due to close-to-zero performance for all models.
These results demonstrate that GPT-family LLMs can decently generate PDDL, while open-source models even up to 405B parameters struggle. As formalizer, gpt-4o-mini, gpt-4o, o3-mini demonstrate non-trivial and increasing performance on BlocksWorld-100. On the more complex Logistics-100, gpt-4o-mini succumbs to zero performance whereas the other two show decreased performance. The solvability of gpt-4o-mini is often much higher than its correctness, suggesting a good grasp of the PDDL syntax but a lack of semantic understanding. In contrast, the solvability of gpt-4o and o3-mini is often 80% to 100% of their correctness. On the other hand, open-sourced models can rarely generate PDDL, a low-resource language, despite them being reportedly strong at generating high-resource languages like Python Cassano et al. (2022). All Llama models up to 405B cannot generate any solvable PDDL across all three datasets, while gemma models show poor though non-zero performance on BlocksWorld-100 and Logistics-100, and strong performance on MysteryBlocksWorld-100.
@@ 5.2 Should LLMs formalize?
Between LLM-as-planner and LLM-as-formalizer, which is the preferred methodology? Figure 3 shows that on BlocksWorld-100, gpt-4o is able to generate solvable PDDL 64/100 times, and of those 64 plans, 60 of them are correct. This far surpasses the LLM-as-planner baseline, which only found correct plans 33/100 times. This trend holds for Logistics-100 as well as the Moderately Templated and Natural BlocksWorld-100 data (Figure 4). On MysteryBlocksWorld-100, we can see that LLM-as-formalizer can generate 70/100 correct plans, which far surpassed LLM-as-planner which did not find a single correct plan as the description becomes unorthodox. The superiority of LLM-as-formalizer also extends to gpt-4o-mini but not o3-mini, who shows strong performance as a planner. These results demonstrate that LLM-as-formalizer greatly outperforms LLM-as-planner in most cases, whenever these LLMs can foramlize PDDL at all. However, these results also show that models that cannot formalize (e.g., Llama models) can still plan, though with close-to-zero performance.
@@ 5.3 The more natural, the harder?
We now examine whether using humanized descriptions makes the problem more difficult. Results from Figure 4 show that on BlocksWorld-100 as the problem sounds more similar to PDDL and less natural, the performance of all the models improves. Similar results hold for the Logistics domain (see results in Appendix C). This suggests that a more natural-sounding domain and problem description is much more challenging than templated, less natural sounding descriptions. One potential explanation is that pattern matching a template back to PDDL is much easier than having to first parse all the predicates and objects from a passage. Another reason is a more natural sounding description may leave out implicit common-sense. For example, the Natural BlocksWorld-100 does not explicitly specify that a block is ‘clear’, because any human who reads that a block is “on top of a stack” can understand that there is no block on top of it and hence ‘clear’ to be moved. However, models often fail to invoke this knowledge and will leave out the ‘clear’ predicate, leading to unsolvable PDDL or incorrect plans.
@@ 5.4 Do LLMs memorize pretraining?
Do LLMs generate plans or formalize PDDL based on what they have memorized in their training data? We determine this by looking at the results on MysteryBlocksWorld-100, a derivative of BlocksWorld where all names are perturbed and nonsensical. From Figure 3, we can see that LLM-as-planner was not able to find a single correct plan using either gpt-4o-mini or gpt-4o. However, gpt-4o-as-formalizer surpassed this baseline with a Correctness score of 70/100. This suggests that LLM-as-formalizer is robust to lexical perturbation, and its success is not due to memorization of the domain which is a part of the pretraining data.
@@ 5.5 What kind of errors?
In this section, we discuss the kind of errors in PDDL generation. We perform an error analysis on a random 20 sample subset of problems where a plan was not found, or the found plan was not correct. From there, we categorize the errors by syntax errors in either file, semantic errors in the $\mathbb{DF}$, and in the $\mathbb{PF}$. Of the errors in the $\mathbb{DF}$, we determine finer-grained errors such as incorrect action preconditions and effects, incorrect or missing predicates, and missing or incorrect action parameters. The error analysis for Natural BlocksWorld-100, Logistics-100 and Heavily Templated MysteryBlocksWorld-100 and can be found in Table 1.

### v3

@@ 4.3 Models
For both of the LLM-as-planner and LLM-as-formalizer approaches, we consider a number of models, including open-source and closed-source LLMs varying in size, including gemma-2-9b|27b-it Team et al. (2024), llama-3.1-8B|70B|405B-Instruct Dubey et al. (2024), DeepSeek-R1-Distill-Llama-8B|70B|671B Guo et al. (2025), gpt-4o-mini-2024-07-18, gpt-4o-2024-08-06, and o3-mini-2025-01-31444platform.openai.com/docs/models. We query these models using KANI Zhu et al. (2023) with default hyper-parameters. The open-source models are run using 4 RTX A6000 GPUs, averaging about 1062 input and output tokens for the LLM-as-formalizer approach in BlocksWorld-100.
To emulate real-life application with minimal user interference, we use zero-shot prompts for all naturalness levels across all datasets (see prompts in Appendix B). We would also like to note that since the LLM has to generate the entire $\mathbb{DF}$ and $\mathbb{PF}$, using few-shot prompting would not be meaningful. Each dataset corresponds to one unique domain and we cannot provide the LLM the $\mathbb{DF}$ in the prompt as the model is supposed to predict it. As for the $\mathbb{PF}$, we also cannot provide it to the model because the typing, predicates, and usage in the $\mathbb{PF}$ would also give away critical information from the file. Nevertheless, we performed a small set of experiments using different prompting methods, such as few-shot and chain-of-thought prompting on Heavily Templated and Natural BlocksWorld-100. Results and discussion of these experiments can be found in the Appendix Sections D.5 and D.6.
@@ 5.1 Can LLMs formalize?
We seek to understand the extent to which LLMs can act as a formalizer to generate entire PDDL, rather than partial components seen in previous work. Figure 3 displays the results of our experiments on Heavily Templated BlocksWorld-100, Logistics-100 and MysteryBlocksWorld-100. Results on the most complex domain Barman-100 is omitted due to close-to-zero performance for all models.
These results demonstrate that GPT-family LLMs and DeepSeek-R1 can decently generate PDDL, while open-source models like Gemma and Llama (even up to 405B parameters) struggle. As a formalizer, gpt-4o-mini, gpt-4o and o3-mini demonstrate non-trivial and increasing performance on BlocksWorld-100. On the more complex Logistics-100, gpt-4o-mini succumbs to zero performance, whereas the other two show decreased performance. The solvability of gpt-4o-mini is often much higher than its correctness, suggesting a good grasp of the PDDL syntax but a lack of semantic understanding. In contrast, the solvability of gpt-4o and o3-mini is often 80% to 100% of their correctness. On the other hand, open-sourced models can rarely generate PDDL, a low-resource language, despite them being reportedly strong at generating high-resource languages like Python Cassano et al. (2022). All Llama models up to 405B cannot generate any solvable PDDL across all three datasets, while Gemma models show poor, though non-zero, performance on BlocksWorld-100 and Logistics-100, and strong performance on MysteryBlocksWorld-100.
@@ 5.2 Should LLMs formalize?
Between LLM-as-planner and LLM-as-formalizer, which is the preferred methodology? Figure 3 shows that on BlocksWorld-100, gpt-4o is able to generate solvable PDDL 64 times, out of 100, and of those 64 plans, 60 of them are correct. This far surpasses the LLM-as-planner baseline, which only found correct plans 33/100 times. This trend holds for Logistics-100 as well as the Moderately Templated and Natural BlocksWorld-100 data (Figure 4). On MysteryBlocksWorld-100, we can see that LLM-as-formalizer can generate 70/100 correct plans, which far surpassed LLM-as-planner which did not find a single correct plan as the description becomes unorthodox. The superiority of LLM-as-formalizer also extends to gpt-4o-mini but not o3-mini, who shows strong performance as a planner. These results demonstrate that LLM-as-formalizer greatly outperforms LLM-as-planner in most cases, whenever these LLMs can formalize PDDL at all. However, these results also show that models that cannot formalize (e.g., Llama models) can still plan, though with close-to-zero performance.
@@ 5.3 The more natural, the harder?
Figure 4: Performance of LLM-as-planner (as-P) and LLM-as-formalizer (as-F) across different naturalness level of description on BlocksWorld-100 and Logistics-100. Detailed results are shown in Appendix D.
We now examine whether using humanized descriptions makes the problem more difficult. Results from Figure 4 show that on BlocksWorld-100 as the problem sounds more similar to PDDL and less natural, the performance of all the models improves. Similar results hold for the Logistics domain (see results in Appendix D). This suggests that a more natural-sounding domain and problem description is much more challenging than templated, less natural sounding descriptions. One potential explanation is that pattern matching a template back to PDDL is much easier than having to first parse all the predicates and objects from a passage. Another reason is a more natural sounding description may leave out implicit common-sense. For example, the Natural BlocksWorld-100 data does not explicitly specify that a block is ‘clear’, because any human who reads that a block is “on top of a stack” can understand that there is no block on top of it and hence ‘clear’ to be moved. However, models often fail to invoke this knowledge and will leave out the ‘clear’ predicate, leading to unsolvable PDDL or incorrect plans.
@@ 5.4 Do LLMs memorize pretraining?
Do LLMs generate plans or formalize PDDL based on what they have memorized in their training data? We determine this by looking at the results on MysteryBlocksWorld-100, a derivative of BlocksWorld where all names are perturbed and nonsensical. From Figure 3, we can see that LLM-as-planner was not able to find a single correct plan using either gpt-4o-mini or gpt-4o. However, gpt-4o-as-formalizer surpassed this baseline with a Correctness score of 70/100. Surprisingly, Gemma models performed very well on the MysteryBlocksWorld-100 dataset with near perfect solvability and correctness for both models. We conclude that the results are mixed across models, which contradicts previous works cited involving PDDL generation which seem to claim that LLM-as-formalizer is clearly better than other methodologies. Despite the inconclusiveness among models, there are more clear-cut results between two groups of models: those that can formalize (eg. gpt-4o and o3-mini) and those that cannot formalize (eg. all Llama models and gpt-4o-mini). Among the models that can formalize, we find that the results are robust to lexical perturbations.
@@ 5.5 What kind of errors?
In this section, we discuss the kinds of errors made in PDDL generation. We perform an error analysis on a random 20 sample subset of problems where a plan was not found, or the found plan was not correct. From there, we categorize the errors by syntax errors in either file, semantic errors in the $\mathbb{DF}$, and semantic errors in the $\mathbb{PF}$. Of the errors in the $\mathbb{DF}$, we determine finer-grained errors such as incorrect or missing actions, preconditions, effects, predicates, and parameters. The error analysis for both Natural BlocksWorld-100 and Logistics-100 can be found in Figure 5.